OpenStudy (anonymous):

The price of a minor league baseball ticket increased and decreased these amounts over four years: -10.8%, +14.2%, +16.5%, -4.7%. What is the MEAN increase or decrease over the four years? (Hint: add and divided by four.)

OpenStudy (shane_b):

The question states how to get the answer: "add and divided by four". What do you get when you do that?

OpenStudy (anonymous):

18.725?

OpenStudy (shane_b):

No...here's what you need to do: \[\frac{(-10.8) + (14.2) + (16.5) + (-4.7)}{4} = ?\]

OpenStudy (anonymous):

i dont get it ;c

OpenStudy (shane_b):

What part don't you understand? The problem is asking you to take the average of 4 values.

OpenStudy (shane_b):

Do you understand how to calculate averages?

OpenStudy (anonymous):

nope

OpenStudy (shane_b):

\[Average = \frac{sum~of~all~values}{number~of~values}\]